A string of length 1.0 m vibrates in 10 loops. If the total mass of string the is 1.0 x 10^ (-3) kg and the tension in it is 10 n, calculate the frequency of the vibration
1
Expert's answer
2020-01-27T10:00:39-0500

The frequency of the string can be computed via the following formula: "f = \frac{1}{2L}\sqrt{\frac{T}{m/L}}", where "L" is the length of the string, "T" is the tension of the string, "m" is the mass of the string.

Substituting "L = 1 m, m = 10^{-3} kg, T = 10 N", obtain "f = 50 {Hz}".